OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 20 = x2 − x?

OpenStudy (anonymous):

A (−4, 0) and (−5, 0) B (−4, 0) and (5, 0) C (4, 0) and (−5, 0) D (4, 0) and (5, 0)

OpenStudy (callisto):

y + 20 = x^2 − x y = x^2 - x - 20 To find x-intercepts, put y=0 into the equation. x^2 - x - 20 = 0 x^2 - 5x +4x - 20 = 0 x (x-5) + 4 (x-5) = 0 (x+4)(x-5) = 0 x+4 = 0 or x-5 = 0 => solve these two equations to get the answers!

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

Welcome :)